Solve using the quadratic formula:
3x2 – 2x = 15x – 10

Answer by stanbon(75887)

Solve using the quadratic formula:
3x2 – 2x = 15x – 10
-----------------------
3x^2-17x+10=0
x = [17 +- sqrt(17^2 - 4*3*10)]/6
x = [17 +- sqrt(169)]/6
x = [17 + 13]/6 or x = [17-13]/6
x = 5 or x = 2/3
================
Cheers,
Stan H.